OpenStudy (anonymous):

Kathy has $1.00 in change. She has at least one of each coin less than half a dollar but does not have half a dollar. What is the least number of coins she has?

ganeshie8 (ganeshie8):

what coins are available in US currency that are less than half dollar ?

OpenStudy (anonymous):

Nickel pennies dimes and quarters @ganeshie8

ganeshie8 (ganeshie8):

may i know how much they each value ?

ganeshie8 (ganeshie8):

1 dollar = how many nickels 1 dollar = how many pennies etc...

ganeshie8 (ganeshie8):

once we have that information, we can setup an equation and solve

OpenStudy (anonymous):

1 dollar= 20 nickels 1 dollar = 10 dimes 1 dollar= 100 pennies 1 dollar = 4 pennies

OpenStudy (anonymous):

I mean 4 quarters*

ganeshie8 (ganeshie8):

lets use this :) http://www.immihelp.com/newcomer/usa-currency-coins.html

ganeshie8 (ganeshie8):

convert everything into pennies, we can setup below linear diophantine equation : \[\large 1p + 5n + 10d + 25q = 100\]

ganeshie8 (ganeshie8):

see if that looks fine ^^

OpenStudy (anonymous):

I'm a confused :/

ganeshie8 (ganeshie8):

since each nickel is worth 5 pennies, having \(n\) nickels is same as having \(5n\) pennies, right ?

ganeshie8 (ganeshie8):

similarly, having \(d\) dimes is same as having \(10d\) pennies, and having \(q\) quarters is same as having \(25q\) pennies

OpenStudy (anonymous):

Yes I get it now :)

ganeshie8 (ganeshie8):

good :) lets see if we can figure out something from the equation : \[\large 1p + 5n + 10d + 25q = 100 \]

ganeshie8 (ganeshie8):

btw, what are we trying to find ?

ganeshie8 (ganeshie8):

this : ` What is the least number of coins she has?`

OpenStudy (anonymous):

Correct

ganeshie8 (ganeshie8):

Clearly she cannot have `4` quarters (why ?)

OpenStudy (anonymous):

@ganeshie8 u r a pro sir

ganeshie8 (ganeshie8):

so hope that she has 3 quarters, and lets find the possible dimes/nickels/pennies

OpenStudy (anonymous):

She has at least one of each coin

ganeshie8 (ganeshie8):

exactly ! so the maximum possible number of quarters is 3 only

OpenStudy (anonymous):

1 dime 2 nickels 5 pennies

ganeshie8 (ganeshie8):

that simplifies the equation to : \[\large 1p + 5n + 10d = 25\]

ganeshie8 (ganeshie8):

again, she cannot have 2 dimes for the same reason, so the equation simplifies to : q = 3 d = 1 \[\large 1p + 5n = 15\]

ganeshie8 (ganeshie8):

You're right ! 3 quarters, 1 dime, 2 nickels, 5 pennies is the solution